mathslover (mathslover):

A tennis ball is thrown vertically upward with velocity of 49 m/s . Calculate 1) the time taken by the ball to reach the max. height . 2) max. height reached by the ball. 3) time taken by the ball to hit the ground after it starts its downward motion from the max. height attained 4) total time taken by it to return to the ground

mathslover (mathslover):

I do found the 1st one ; Time of ascent = \(\large{\frac{u sin \theta}{g}}\) = 49 / 9.8 = 490 / 98 = 5 seconds m i right?

mathslover (mathslover):

or we can do like this : v = u -gt 0 = 49 - (9.8) T -49/- 9.8 = T 5 seconds = Time of ascent

OpenStudy (kinggeorge):

That looks right to me.

mathslover (mathslover):

ok now coming to the second part: H = \(\large{\frac{u^2 sin^2 \theta}{2g}}\) 49^2 / 2*9.8 = 49 * 49 / 19.6 = 122.5 m

mathslover (mathslover):

right?

OpenStudy (kinggeorge):

Looks great.

mathslover (mathslover):

i am stuck at this point now :time of descent

OpenStudy (kinggeorge):

Since acceleration is the same, it must be the same amount of time it took to reach the max height. Thus, the descent time is 5 seconds.

mathslover (mathslover):

well right but according to the formula : tiem fo descent = u sin theta / g

mathslover (mathslover):

but when : the ball comes down the final velocity will become initial velocity there

OpenStudy (kinggeorge):

right...

mathslover (mathslover):

so this is ..

mathslover (mathslover):

we have : v1 = 0 m/s , u1 = 49 m/s v1 = u2 = 0 m/s and u1 = v2 = 49 m/s

mathslover (mathslover):

but in the formula : T = u sin theta / g we have u not v that is initial velocity 2 : u2 but u2 = 0 *u2 does not mean u^2 it means u_2 hence T = 0

mathslover (mathslover):

where m i wrong? kg?

OpenStudy (kinggeorge):

In this equation, \(u\) is used as the initial velocity? Or the velocity at ground?

mathslover (mathslover):

I think u is used as the initial velocity wait lemme confirm it

mathslover (mathslover):

i know the deriviation of that formula though .. may be that help? : \[\large{s=ut -\frac{1}{2}gt^2}\] \[\large{0 = usin\theta t-\frac{1}{2}gt^2}\] \[\large{t = 0 , t = \frac{2usin\theta}{g}}\]

OpenStudy (kinggeorge):

The first equation is not correct if you are starting at a non-zero height. The full equation is\[\large{s=s_0+ut -\frac{1}{2}gt^2}\]

mathslover (mathslover):